QD = 25 - 3P and QS = 10 + 2P, write the inverse demand function


1
Expert's answer
2022-01-27T13:15:41-0500

"Q_D= 25-3P"

Inverse demand function= Expressing P as a function of Q

"\therefore 3P= 25- Q_D"

"P=\frac{25-Q_D}{3}"